OpenStudy (anonymous):

ONE QUESTION HELP PLEASE! how is a graph suppose to look like with this equation? correctly represents the equation 3x + 4y = 12.

OpenStudy (anonymous):

and how does aline grapgh look like to this problem: Which graph represents the solutions to the inequality |2x − 6| < 4?

OpenStudy (anonymous):

What is the equation of the line that passes through the points (−2, 1) and (1, 10)? 3x − y = −7 3x − y = −5 3x − y = 5 x + 3y = −5

OpenStudy (anonymous):

anybody i tihnk my wi fi is broken!?

OpenStudy (lasttccasey):

For your first question see the attached picture and look at this for reference: http://www.wolframalpha.com/input/?i=3x+%2B+4y+%3D+12

OpenStudy (lasttccasey):

#2, Look at the attached picture and check out this link for reference, http://www.wolframalpha.com/input/?i=%7C2x+%E2%88%92+6%7C+%3C+4

OpenStudy (lasttccasey):

For the last one, http://www.wolframalpha.com/input/?i=equation+of+a+line++%28%E2%88%922%2C+1%29+and+%281%2C+10%29 Answer is 3x − y = −7

OpenStudy (lasttccasey):

Can be rewritten as,\[y=3x+7\]